OpenStudy (anonymous):

what is the reference angle of -9pi/7?

OpenStudy (anonymous):

i got that it was 51 degrees, but i need to put that into radians

OpenStudy (anonymous):

and im not quite sure how^

OpenStudy (unknownunknown):

It's already in radians.

OpenStudy (anonymous):

i need to put it in terms of pi.

OpenStudy (anonymous):

In general,\[\pi = 180\] so lets do the calculation: \[\frac{ -9 \times 180 }{ 7 } \approx -231\]

OpenStudy (unknownunknown):

-73.666

OpenStudy (anonymous):

okay i that makes sense. but i need my answer to be in fraction form

OpenStudy (anonymous):

here are my option choices.

OpenStudy (anonymous):

so then -231,428571 degrees

OpenStudy (anonymous):

yes..do you see my answer choices there? how is it that i put it into that form?